# 00footnotetext:  $\ast$ Corresponding author

### v1

@@ Abstract
  • We also give a lower bound estimate (in terms of the length of the longest side) for the Type-A first mixed eigenvalue of triangles using another approach. This lower bound in some cases is more sharper.
@@ 1 Introduction
Let $T\subset\mathbb{R}^{2}$ be a planar triangle in the Euclidean plane $\mathbb{R}^{2}$, and let $L$, $M$, $S$ be the lengths of sides of the triangle $T$ such that $L\geq M\geq S$. Denote by $\lambda_{1}^{set}$ the smallest eigenvalue corresponding to the Dirichlet conditions applied to a chosen set of sides – in this setting, e.g., $\lambda_{1}^{LS}$ would correspond to the Dirichlet conditions imposed on the longest and shortest sides. Consider the mixed Dirichlet-Neumann eigenvalue problem of the Laplacian on the triangle $T$ as follows
@@
$\lambda_{1}^{MS}(T)\geq\frac{4\pi^{2}}{\ell^{2}},$ (1.5)
where $\ell$ is the length of the longest side $L$ of the triangle $T$.
@@
We wish to make a comparison between the estimate (1.4) and the estimate (1.5). Let $\alpha$ be the angle corresponding to the longest side $L$, and let $m$, $s$, $\ell$ be the lengths of sides $S$, $M$ and $L$ of the triangle $T$, respectively. We divide the comparison into two cases:
Case I. When $\alpha\geq\frac{\pi}{2}$, one has
@@
This is to say that when $\alpha\geq\frac{\pi}{2}$ (i.e. $T$ is a right triangle or an obtuse triangle), the lower bound estimate (1.4) is sharper than (1.5). Especially, if $\ell^{2}=4|T|$, then $m=s$ and $\alpha=\pi/4$, which implies that in this setting $T$ should be an isosceles right triangle.
@@
that is, $\ell^{2}<4|T|$. So, in this setting, the lower bound estimate (1.5) is sharper than (1.4).
@@
which means $4|T|>\ell^{2}$ in this setting; if $m=4$, $s=2$, a simple calculation tells us
$4|T|\approx 15.454,\qquad\ell^{2}\approx 17.93,$
@@
In sum, through the above comparison, readers can see that it is meaningful to give estimates (1.4)-(1.5) for the mixed eigenvalue $\lambda^{MS}_{1}(T)$ simultaneously, and these estimates deeply and intuitively reveal the relation between the eigenvalue $\lambda^{MS}_{1}(T)$ and the length of the longest side, the area of the given triangle $T$.
@@ 2 The case of triangles
$\phi(x)=\left\{\begin{array}[]{llll}v(x)&\mathrm{if}~x\in T_{R},\\ v(y)&\mathrm{if}~x\in\triangle BCD,\\ v(z)&\mathrm{if}~x\in\triangle ACE,\\ v(w)&\mathrm{if}~x\in\triangle DCE,\end{array}\right$.
@@
$\phi=\left\{\begin{array}[]{lll}u&\mathrm{on}~~T,\\ u_{13}&\mathrm{on}~~\triangle GAB,\\ u_{24}&\mathrm{on}~~\triangle AHC,\end{array}\right$.
@@
where $\mu_{2}(\mathcal{C})$ is the nonzero Neumann eigenvalue of $\mathcal{C}$. By using the basic properties of triangles, one has $0<h\leq\sqrt{3}\ell/2$, and moreover $h=\sqrt{3}\ell/2$ if and only if $T$ is an equilateral triangle. This fact, together with (2.7), implies
$\mu_{2}(\mathcal{C})=\frac{4\pi^{2}}{\ell^{2}}.$
Due to the definition of $\phi$, it is not hard to see $\int_{\mathcal{C}}\phi=0$, which implies that $\phi$ can be regarded as a trial function for $\mu_{2}(\mathcal{C})$. Hence, we can obtain
@@
$\geq$ $\mu_{2}(\mathcal{C})=\frac{4\pi^{2}}{\ell^{2}},$
which completes the proof of Theorem 1.7. ∎
@@ 3 The case of trapezoids
$\lambda_{k}^{Q^{1}Q^{2}}(\Omega)\int_{\Omega}f^{2}\leq\int_{\Omega}|\nabla f|^{2}=\sum\limits_{j=1}^{k}\alpha_{j}^{2}v_{j}\leq v_{k}\int_{\Omega}f^{2},$
@@
$z_{+}:=\frac{z+|z|}{2}=\left\{\begin{array}[]{ll}z,&when~z\geq 0\\ 0,&when~z<0\end{array}\right$.
@@
$\pi^{2}\left(\frac{i^{2}}{m^{2}}+\frac{h^{2}}{b^{2}}\right),\qquad i=1,2,3,\cdots,~~j=0,1,2,\cdots.$
@@
$\pi^{2}\left(\frac{i^{2}}{m^{2}}+\frac{h^{2}}{b^{2}}\right)=\frac{\pi^{2}}{m^{2}}\left(i^{2}+j^{2}\cdot\frac{m^{2}}{h^{2}}\right)>\frac{\pi^{2}}{m^{2}}\left[i^{2}+j^{2}\cdot(k^{2}-1)\right].$
@@
$\rho(x)=\left\{\begin{array}[]{ll}v(x),&\mathrm{if}~x\in\Gamma_{R},\\ v(y),&\mathrm{if}~x\in I_{s}\setminus\Gamma_{R},\end{array}\right$.
@@ Acknowledgments
This research was supported in part by the NSF of China (Grant Nos. 11801496 and 11926352), the Fok Ying-Tung Education Foundation (China), Hubei Key Laboratory of Applied Mathematics (Hubei University), and Key Laboratory of Intelligent Sensing System and Security (Hubei University), Ministry of Education.

### v2

@@ Abstract
  • We also give a lower bound estimate (in terms of the length of the longest side or the length of the height of the longest side) for the Type-A first mixed eigenvalue of triangles using another approach.
@@ 1 Introduction
Let $T\subset\mathbb{R}^{2}$ be a planar triangle in the Euclidean plane $\mathbb{R}^{2}$, and let $L$, $M$, $S$ be the sides of the triangle $T$ such that their lengths satisfy $\mathcal{L}(L)\geq\mathcal{L}(M)\geq\mathcal{L}(S)$. Denote by $\lambda_{1}^{set}$ the smallest eigenvalue corresponding to the Dirichlet conditions applied to a chosen set of sides – in this setting, e.g., $\lambda_{1}^{LS}$ would correspond to the Dirichlet conditions imposed on the longest and shortest sides. Consider the mixed Dirichlet-Neumann eigenvalue problem of the Laplacian on the triangle $T$ as follows
@@
$\lambda_{1}^{MS}(T)\geq\min\left\{\frac{\pi^{2}}{h^{2}},\frac{4\pi^{2}}{\ell^{2}}\right\},$ (1.5)
where $\ell$ is the length of the longest side $L$ of the triangle $T$, and $h$ is the length of the height of the side $L$.
@@
Let $\alpha$ be the angle corresponding to the longest side $L$, and here let $s$, $m$, $\ell$ be the lengths of sides $S$, $M$ and $L$ of the triangle $T$, respectively. We divide our argument into two cases:
Case I. When $\alpha\geq\frac{\pi}{2}$ (i.e. $T$ is a right triangle or an obtuse triangle), one has
@@
Especially, if $\ell^{2}=4|T|$, then $m=s$ and $\alpha=\pi/2$, which implies that in this setting $T$ should be an isosceles right triangle.
@@
that is, $\ell^{2}<4|T|$.
@@
which means $4|T|>\ell^{2}$ in this setting. However, if $\ell=4$, $m=3.9823$, $s=2$, a simple calculation tells us
$4|T|\approx 15.4548,\qquad\ell^{2}=16,$
@@
In sum, we have:
  • (a) If $T$ is an obtuse triangle, a right triangle (except the case of isosceles right triangles), or an acute triangle which has the form like the triangle $\ell=4$, $m=3.9823$, $s=2$ constructed in (2) of Case II, then $\ell^{2}>4|T|$.
  • (b) If $T$ is an isosceles right triangle, then $\ell^{2}=4|T|$.
  • (c) If $T$ is an isosceles acute triangle or an acute triangle which has the form like the triangle $m=4$, $s=3$, $\alpha=5\pi/12$ constructed in (2) of Case II, then $\ell^{2}<4|T|$.
On the other hand, it is not hard to see that:
  • If $0<h<\frac{\ell}{2}$, then $h^{2}<\frac{\ell^{2}}{4}$ and $|T|<\frac{\ell^{2}}{4}$. In this situation, $T$ must be a triangle belonging to the category (a) described as above, and moreover, the estimate (1.5) in Theorem 1.7 becomes $\lambda_{1}^{MS}(T)\geq\frac{4\pi^{2}}{\ell^{2}}$. Since $|T|=\frac{h\ell}{2}<\frac{\ell^{2}}{4}$, one knows that in this situation, the estimate (1.4) is sharper than the estimate (1.5).
  • If $h=\frac{\ell}{2}$, then $h^{2}=\frac{\ell^{2}}{4}$ and $|T|=\frac{\ell^{2}}{4}$. In this situation, $T$ must be an isosceles right triangle, i.e. belonging to the category (b). Moreover, the estimate (1.5) becomes $\lambda_{1}^{MS}(T)\geq\frac{\pi^{2}}{|T|}$, which coincides with the lower bound estimate given in (1.4).
  • If $h>\frac{\ell}{2}$, then $h^{2}>\frac{\ell^{2}}{4}$ and $|T|>\frac{\ell^{2}}{4}$. In this situation, $T$ must be a triangle belonging to the category (c), and moreover, the estimate (1.5) becomes $\lambda_{1}^{MS}(T)\geq\frac{\pi^{2}}{h^{2}}$. Since $|T|=\frac{h\ell}{2}<h^{2}$, one knows that in this situation, the estimate (1.4) is sharper than the estimate (1.5).
Therefore, if $T$ is an isosceles right triangle, then the lower bound in the estimate (1.4) or (1.5) coincides with each other, while if $T$ is an obtuse triangle, a right triangle (except the case of isosceles right triangles), or an acute triangle, then the lower bound estimate (1.4) is strictly sharper than the estimate (1.5).
@@ 2 The case of triangles
$\phi(x)=\left\{\begin{array}[]{llll}v(x)\qquad&\mathrm{if}~x\in T_{R},\\ v(y)\qquad&\mathrm{if}~x\in\triangle BCD,\\ v(z)\qquad&\mathrm{if}~x\in\triangle ACE,\\ v(w)\qquad&\mathrm{if}~x\in\triangle DCE,\end{array}\right$.
@@
$\phi=\left\{\begin{array}[]{lll}u\qquad&\mathrm{on}~~T,\\ u_{13}\qquad&\mathrm{on}~~\triangle GAB,\\ u_{24}\qquad&\mathrm{on}~~\triangle AHC,\end{array}\right$.
@@
where $\mu_{2}(\mathcal{C})$ is the nonzero Neumann eigenvalue of $\mathcal{C}$. Due to the definition of $\phi$, it is not hard to see $\int_{\mathcal{C}}\phi=0$, which implies that $\phi$ can be regarded as a trial function for $\mu_{2}(\mathcal{C})$. Hence, we can obtain
@@
$\geq$ $\mu_{2}(\mathcal{C}),$
which, together with (2.7), completes the proof of Theorem 1.7. ∎
@@ 3 The case of trapezoids
$\lambda_{k}^{Q^{1}Q^{2}}(\Omega)\int_{\Omega}f^{2}\leq\int_{\Omega}|\nabla f|^{2}=\sum\limits_{j=1}^{k}\alpha_{j}^{2}v_{j}\int_{\Omega}f^{2}\leq v_{k}\int_{\Omega}f^{2},$
@@
$z_{+}:=\frac{z+|z|}{2}=\left\{\begin{array}[]{ll}z,\quad&when~z\geq 0\\ 0,\quad&when~z<0\end{array}\right$.
@@
$\pi^{2}\left(\frac{i^{2}}{m^{2}}+\frac{j^{2}}{h^{2}}\right),\qquad i=1,2,3,\cdots,~~j=0,1,2,\cdots.$
@@
$\pi^{2}\left(\frac{i^{2}}{m^{2}}+\frac{j^{2}}{h^{2}}\right)=\frac{\pi^{2}}{m^{2}}\left(i^{2}+j^{2}\cdot\frac{m^{2}}{h^{2}}\right)>\frac{\pi^{2}}{m^{2}}\left[i^{2}+j^{2}\cdot(k^{2}-1)\right].$
@@
$\rho(x)=\left\{\begin{array}[]{ll}v(x),\qquad&\mathrm{if}~x\in\Gamma_{R},\\ v(y),\qquad&\mathrm{if}~x\in I_{s}\setminus\Gamma_{R},\end{array}\right$.
@@ Acknowledgments
This research was supported in part by the NSF of China (Grant Nos. 11801496 and 11926352), the Fok Ying-Tung Education Foundation (China), the Key Project of Jiangxi Provincial Natural Science Foundation (Grant No. 20252BAC250003), Hubei Key Laboratory of Applied Mathematics (Hubei University), and Key Laboratory of Intelligent Sensing System and Security (Hubei University), Ministry of Education.
Conflict of interest
The authors declare that there are no conflicts of interests regarding the publication of this paper.
Data availability statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.